OpenStudy (anonymous):

How do you solve this problem? 5x-4=square root 5x-4?

OpenStudy (anonymous):

By eye ball, x=1 will work if\[5x-4=\sqrt{5x-4} \]

OpenStudy (anonymous):

Please I need to see the step to learn.

OpenStudy (anonymous):

Square and expand each side.\[(5x-4)^2=\left(\sqrt{5x-4}\right)^2 \]\[25 x^2-40 x+16=5 x-4 \]Move the RHS to the LHS and simplify.\[16-40 x+25 x^2-(-4+5 x)=0\]\[25 x^2-45 x+20=0 \]The LHS factors as follows:\[5 (x-1) (5 x-4)=0 \]Can you complete the solution from here?

OpenStudy (anonymous):

next step please?

OpenStudy (anonymous):

x=5 and x=-4

OpenStudy (anonymous):

(x-1) = 0 x = 1 (5x-4) = 0 x = 4/5

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

You're welcome.